OpenStudy (anonymous):

A jumbo jet starting from rest needs to attain a speed of 350 km/hr on the runway before it can take off. Can it use a runway 1km long if it is to accelerate at the rate of 2.67m/s squared? please explain why

OpenStudy (kropot72):

First convert the speed of 350 km/hr to the same units as the acceleration: \[\large 350\ km/hr=\frac{350\times1000}{3600}=97.22\ m/s\] The time to accelerate from rest to 97.22 m/s is given by dividing the maximum speed by the acceleration as follows: \[\large t=\frac{97.22}{2.67}=36.41\ seconds\] The average speed of the aircraft is 97.22/2 = 48.61 m/s. The distance traveled at an average speed of 48.61 m/s in 36.41 s is given by: \[\large48.61\times36.41=1769.89\ meters\] Will a runway only 1000 meters long be adequate?